a uniform beam of mass M = 61 kg and length L = 11 m is hung horizontally by vertical cables at the two ends, as shown. a mass of m = 31 kg is then hung from the beam at a distance of 5/8L from the left side. what is the tension in the rightmost cable?

wHd+AKZsmRNCwAAAABJRU5ErkJggg==
Answer #1

There are two type of equilibrium here , one is translational and other one is rotational

For translational , let the tensions be T1 and T2 , T2 being the tension in Rightmost cable

T1 + T2 = 61g + 31g = 92g

Also for rotational equilibrium

T1 *L/2 - 31g*(5/8L-L/2) -T2*(L/2) =0 ( taking torque about centre of the cable

T1/2 -31g*1/8 -T2/2 = 0

T1-T2 = 31g/4

Substraction of second from first

2T2 = 92g-31g/4

T2 = 42.125g = 412.825N